the sum of two numbers is 42. One number plus two times the other number is 57. What are the numbers? The answers are 15 and 27 but I don't understand how to get there. Could you help?
Answer by venugopalramana(3286)

the sum of two numbers is 42. One number plus two times the other number is 57. What are the numbers? The answers are 15 and 27 but I don't understand
let one number be =x
let the other number be =y
their sum =x+y=42..........(1)
two times the other number =2y
plus the first number x = x+2y=57...........(2)
(2)-(1) ....x+2y-x-y=57-42=15
y=15...substituting in (1),we get ..x+15=42
x=42-15=27